SW&IT Meeting Minutes 2018-06-11
Attendees: Andrew, David, Doug, Igor, Joe, Kellee, Matthieu, Petr, Stephen, Xianyu, Yang
News from management
- OSA interview
- PBP meeting with Roberto at SPIE, readiness?
Previous Action Items
- revision control workflows [IS/PK] after SPIE
Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.
- SOUL server update ?
- TCS eng. night results ?
- circulate SSD plan (Leroy and Chris in the loop) [MB]
- git migration, next? [MB]
This week in github
- coding standards, indent tools, etc.
Last week's epidose:
- Finsihed TRAC to Wiki conversion
- minor changes still needed (update obsolete hyperlinks, spelling, etc.)
- Completed first draft of handover document
- ARGOS handover meeting with MPIA
- Send handover document to MPIA
- Get TRAC backup from Florian
- Jenkins configuration
- SVN --> GitHub
- Add MPIA users to Wiki
- ARGOS users cleanup
- Verified IT's claim that they do parse email lists, so we've changed direction and will change all IT distribution lists to IT internal groups/lists
- Working on documenting the current state so we can pass that around and make sure all the lists make sense
- Got Luci Sun machine working for Dave to pull configuration files; he says that he has what he needed.
- Finish up documentation for IT lists
- Create some new lists
- Backup testing
Spend some more time on the Scheduler current view. I'm trying to leave it in a "usable" state, even though is wont be perfect it could still be used. So I'm testing a lot and correcting bugs where possible, other bugs won't be corrected, at least not until the LUCI
software handover. In addition to this the code (the current view) would benefit of a serious design pattern overhaul, but again maybe this is something that can be done after the software handover.
Most of my time was spent assesing some of the ideas put forward in the AFC OT meetings. Looking into the code and partially implementing some of the ideas.
- Planner can now work with n Science Programs concurrently (DONE)
- Focused program is determined by the user selecting a Program node.
- I chose this over other possibilities because the TPE works that way and users will know what to expect. ME agrees.
- Toggle select is working for all cases now (DONE)
- The menu items for adding / removing observations from a plan are multi-select now.
- Make sure that changes to Target on the OT are propagated to the Planner.
- Fix a drawing bug that showed up during Thursday's meeting (ME, IS)
- I am off Thursday + Friday for vacation
- AGW3 WFS camera (on LUCI1) stopped producing images - researched old email, notes, etc. to help with troubleshooting.
- Set up the 3.4.0 UEI software and recompiled new OVMS version using that version - multiple discussions with support to get more info about the new version.
- GCS -- looked into flip for image display. It's not there - Petr will implement it.
- Checked new OAC running for errors/warnings, but looks like they haven't done any testing. And since the AGW3 was having problems they were all pulled in to that task.
- Vacation day on Thursday.
- MODS software handover status meeting with Rick, Matthieu.
- Add MODS pressures to ALH.
- Fix LBC power control of filter wheels in error conditions.
- Work on GCS updates for OAC HK upgrade.
- Support engineering tests of OAC HK upgrade.
- Vacation day on Thursday.
- management OC, WPC
- ARGOS handover planning and meeting with MPIA
- coordinate LUCI Sun back to Tucson
- iLocater in IT
- PEs (continued)
- OSA interview
- obsdev meeting
- PBP meeting
- Ketiv wrap up
- rehost/package ims (issuetrack web tool)
- MODS handover meeting with OSU
- LUCI handover meeting with Alexander @ SPIE (TBC)
- finished fix for co-pointing
- added much more tests for PCS
- hopefully fixed PCS died (race conditions accessing vector)
- tme to limits
- GCS guide cam flip
- installed SOUL computer, necessary plumbing.
- ARGOS meeting
- Big displays repaired after Louise warranty work.
- Clone for OVMS work this summer.
- Need confirmation that LUCI "CCD Connection" box can be removed from rack.
- More work on IT7260 and installed the changes
- Operation support, e.g., AdSec elevation limit change, switch BCU port change, resolve issues for adsec power off by safety, disable/enable safety monitoring script
- Updated the PBP SW plan
- Documented the current actions and new PBP parameters in housekeeper, fastdiagnostic and MOXA
- Support the SOUL testing, e.g., let sxadsec work with SOUL, organize our SOUL box installation, communicate to the mountain for LBTI-SOUL test, and discuss issues with SOUL people for issues
- Continuing the work on IT#7261 (WfsArbitrator memory issues):
- Study the core dumps from last week when memory usage is high (but may not find much useful information).
- Produce a continuous monitoring script to record memory changes.
- Continue to study part of the code base (current focus is still on the Python wrapper library).
- Continue to work on AO document wiki pages.
- Start work on Collimation and Pointing wiki page. Initially it will contain all relevant documents we have and could find. Later time I'll be writing a more direct and cohesive documentation gradually when I gain more understanding of the process.
- Study pointing model related documents.
- Study and experiment on current existing scripts and programs on producing pointing models.